# NeuroPrune: A Neuro-inspired Topological Sparse Training Algorithm for Large Language Models

### v1

@@ 1 Introduction
Refer to captionRefer to caption
Figure 3: Left is the fraction of times a head is removed using NeuroPrune when fine-tuning on SST2 with a BERT model. The overall numbers (blue curve) are averaged across layers and runs ($\pm$ sd), where at least $10$ heads are removed. We also show individual layer numbers averaged across runs for the top three layers where most pruning of heads happens (Figure 10 in appendix). We see there is a significant bias towards keeping the last head in each layer leading to a more modular structure and also showcasing preferential attachment, as neurons from the previous and next layers will connect only to these heads. The middle head (head $5$) is also maintained more than most other heads, as it replaces many of the earlier heads. Results averaged across GLUE tasks on the right are similar.
Redundancy-based pruning While structured sparsity aims at preferential attachment, such pruning (by zeroing out weights) cannot determine which connections are redundant. Elimination of redundant connections is an important aspect of the refinement process Wang et al. (2011); Hashimoto and Kano (2013) that takes place after the brain develops very dense networks of connections. In the case of LLMs, we hypothesize that such redundancy can be measured by similarity between attention heads, whereupon similar attention heads can then be merged, resulting in reduced complexity while maintaining functionality (i.e. performance is maintained on downstream tasks). Such removal of redundancy is conjectured in Lichtman and Colman (2020) to be unique to the neuron development in the central nervous system of vertebrates. Figure 3 illustrates how often heads are found to be redundant using NeuroPrune. Generally, the last head is found to be the least redundant, while the middle head also exhibits limited redundancy. In Figure 10 in the appendix, we see head removal as a function of layers and find that the last three layers have the highest number of redundant heads.
@@
Our approach has the following benefits: 1) It is task agnostic. 2) It is easily adaptable to different transformer-based LLM architectures as it does not add additional (mask) variables to do the pruning. We apply it to BERT (encoder), T5 (encoder-decoder) and OPT (decoder) models. 3) It learns sparsity patterns exhibiting principled topological structure. 4) It results in LLMs with a competitive and even sometimes superior performance on different benchmarks and tasks (GLUE, summarization, machine translation), although our proposal is more neuroscience-motivated than solely trying to maximize performance. 5) It is generally much faster to train than the competing baselines, with time per epoch being similar to standard fine-tuning. It also exhibits inference speedups as the topological constraints encourage $N$:$M$-type sparsity.
@@ 4 Method
Input: Model $M(\cdot;A_{0},L_{0})$ with $N_{L}$ layers, initial Query, Key, Value concatenated matrices $A^{(l)}_{0}$ and MLP weight matrices $L^{(l)}_{0}$ for layer $l$, sparsity parameters $\alpha,\beta\geq 0$, # epochs $n_{e}$, loss function $\lambda$, redundancy threshold $\theta$
@@
Algorithm 1 NeuroPrune (Code provided)
@@ 4.1 MLP Sparsification
$R_{mlp}^{(l)}(L^{(l)})=$ $\frac{1}{d}[n^{(l)}_{in,1},...,n^{(t)}_{in,d_{e}}]\cdot|L_{in}^{(l)}|\cdot\vec{1}_{d}$ (1)
$+\frac{1}{d_{e}}[n^{(l)}_{out,1},...,n^{(t)}_{out,d}]\cdot|L_{out}^{(l)}|\cdot\vec{1}_{d_{e}}$
@@ 4.2 Attention Sparsification
Note that the above constraint is applied across heads in the attention layer as it considers the entire $Q$, $K$, $V$ matrices (hence the inner summation over $3d$ entries). Additionally, while the standard $l_{2}$ group penalty induces weights within a group to be similar, this $l_{1}^{.5}$ group penalty allows sparsity patterns to be learned within a group. For example, in Figures 1 and 7, while entire rows are often removed, we also observe that certain rows only exhibit sparsity in $Q$ and $K$ while leaving corresponding rows of $V$ dense, which is still valuable as it indicates that attention may not be required for those neurons/dimensions.
@@ 4.4 NeuroPrune
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@ 5 Experiments
Metrics: For performance we report accuracy (Acc) for the GLUE datasets except COLA where Mathew’s Correlation (MCorr) is a standard metric, Rouge for summarization and BLEU scores for machine translation. We also report (average) inference and train times.
Experimental details such as the hyper-parameter values explored for each method, the number of epochs they are run, and the hardware used are in the appendix.
@@ 5.1 Varying sparsity percentage
GLUE: In Figure 4 we see the behavior of our method on GLUE datasets for BERT, T5, and OPT models. We see that our method is always competitive with CoFI, outperforming it on the smaller datasets. This is possible because we do not add extra variables to the model, which when coupled with the topological constraints results in stabler performance. The structured sparsity also gives improved inference times and the train time is much lower, since not only do we need to run our method only for a few epochs, but the time per epoch is the same as standard fine-tuning.
For T5 and OPT we compare with $l_{1}$ sparsification. As can be seen, NeuroPrune is better than $l_{1}$ in most cases w.r.t. performance, especially for in-between sparsities. We believe this happens because NeuroPrune can choose between attention or MLP ($\alpha$, $\beta$ parameters), on which to sparsify more when optimizing performance even though the constraints are structured. We also see benefits in inference time, again possibly because of the structured sparsification. Train times are similar as we run $l_{1}$ for the same number of epochs as NeuroPrune and its per-epoch time is similar to that of standard fine-tuning.
@@ 5.2 Varying number of heads
Figure 6: Performance (left) and Efficient Frontier (right) for NeuroPrune and DSP on a German to English translation task. NeuroPrune consistently outperforms DSP across models with varying numbers of attention heads present and typically with much high levels of sparsity. The efficient frontier shows that for similar levels of performance, NeuroPrune finds much sparser models for a fixed number of heads present.
@@ 5.3 Other Insights
Topological sparsity: As seen in Figure 1 (and Figures 7, 8, 9 in the appendix) we see that our constraints try to eliminate neurons both in the MLP layers as well as the attention layers. This is observed as a row/column (sparsity) pattern in these matrices. Interestingly, sometimes only the $Q$ and $K$ entries in a row are sparsified, although the group sparsity constraint is applied to $QKV$ jointly. The similarity in sparsification patterns for matrices $Q$ and $K$ directly reflects the symmetry of the operations they are jointly involved in, when computing the attention coefficients as inner products of each row of $Q$ with each row of $K$, compactly as $QK^{\top}$, unlike $V$ which is simply a linear projection of the token embeddings. The MLP portions, as seen in Figure 1, exhibit preferential attachment with increasing sparsity which is consistent with brain functional networks.
Importance of removed heads: In Figure 11 in the appendix, we see the relative importance of removed heads, defined as the ratio between the importance of removed heads in a layer to the ones that remain averaged across all the layers. The individual head importances are computed as the sum of the absolute output dense weights that emerge from a head. We see that NeuroPrune can prune more important heads than DSP as it finds redundant heads, thereby sparsifying more aggressively, which is evidenced by the faster training time on GLUE tasks for similar levels of head removal.
Head removal: As seen in Figure 5 there is a high bias to keep the last head in each layer. This is because NeuroPrune prefers keeping later heads in a layer if it is similar to a set of heads that another head may be similar to, thus encouraging more modularity in a layer. In Figure 10 in the appendix, we see that later layers, and often the first layer, lose more heads than the intermediate ones, which is consistent with Li et al. (2021).
@@ 6 Discussion
This work shows that NeuroPrune, inspired by sparsity, modularity, and preferential attachment topology in brain functional networks, is competitive with other SOTA dynamic sparse training methods, even though it does not solely try to optimize performance. It is also more efficient than them in train time with speed-ups also seen in inference.
There are multiple avenues for future research. First, it would be interesting to combine our redundancy-based and head-importance pruning methods to produce even more aggressive and efficient pruning of heads. Second, the structured strategies we used for fine-tuning could also be tried during pre-training. Third, one could test the generalizability of models trained using NeuroPrune on related, but different tasks and measure if similar gains can be secured. We believe our work will spur progress on efficient architectures.
@@
All the datasets we considered were for the English language. Results may vary for other languages. We applied our method to three LLMs, but more architectures could be tested in the future not to mention more tasks. Our method although easy to adapt to different architectures while also being efficient, it does not allow the user to specify the exact number of heads one wants to prune and architectures are limited to Transformer based LLMs. This is implicitly a function of the threshold $\theta$ and the similarity of the pre-trained/fine tuned heads as well as the sparsity. We also have three hyper-parameters ($\alpha$, $\beta$ and $\theta$) that need to be specified for each run.
@@
Appendix A Experimental Details
Setup: NeuroPrune results were obtained by varying $\alpha$ and $\beta$ from $10^{-7}$ to $0.1$ in multiples of $10$. The $l_{1}$ penalty parameter also took these values. $\theta$ took values in $\{0.15,0.2,0.25\}$ for the head removal experiments where the default was set to $0.15$ for the GLUE and summarization tasks. It was sufficient to run NeuroPrune for a single epoch for the larger GLUE datasets (viz. MNLI, QQP, QNLI and SST2) and the summarization task, while for the other smaller GLUE datasets we ran it for $3$ to $5$ epochs. $\epsilon$ for MLP sparsification was set to $10^{-4}$. CoFI finetunes the model before it starts pruning. For smaller GLUE datasets, finetuning before pruning epochs is $4$ and total epochs is $100$, while for larger datasets, these numbers are $2$ and $20$ respectively. All the other parameters were unchanged. To make the comparison fair, we turned off the distillation option. DSP runs 3 epochs of joint (finetuning and mask learning) training. For machine translation, NeuroPrune results were obtained with $\alpha\in\{0.005,0.01,0.05,0.1,0.25,0.5\}$, $\beta\in\{0.05,0.1\}$ and $\theta$ varying from 0.1 to 0.4 in increments of 0.02. $\epsilon$ was set as in GLUE. Since there was no pretrained model, 15 epochs of pretraining were done followed by 15 epochs of NeuroPrune finetuning for a total of 30 epochs. DSP joint training was also done with a total of 30 epochs.
@@ Appendix A Experimental Details
Figure 10: We see the (average) # of heads removed per layer using NeuroPrune when fine tuning on the GLUE Benchmark datasets with a BERT(-base) model for cases where at least 10 heads are removed. As can be seen more heads are pruned from the later layers and the first one as compared to the middle layers.
Figure 11: Relative importance of removed heads for NeuroPrune and DSP averaged across runs, where at least 10 heads are removed. We see that NeuroPrune removes more important heads than DSP does because of its redundancy based head elimination algorithm.
Appendix B Additional Figures
In Figures 7, 8 and 9 we sparsity patterns induced by NeuroPrune in a BERT-base model at different sparsity percentages on the SST2 dataset.
In Figure 10, we see the number of heads removed on average per layer over the GLUE benchmark. As can be seen, the last three layers and the first layer undergo most pruning.
In Figure 11, we see the relative head importance of removed to kept heads for NeuroPrune and DSP. As can be seen our redundancy based pruning removes more important heads on average.

### v3

@@ 1 Introduction
Figure 3: Left is the fraction of times a head is removed using NeuroPrune when fine-tuning on SST2 with a BERT model. The overall numbers (blue curve) are averaged across layers and runs ($\pm$ sd), where at least $10$ heads are removed. We also show individual layer numbers averaged across runs for the top three layers where most pruning of heads happens (Figure 7(right)). We see there is a significant bias towards keeping the last head in each layer leading to a more modular structure and also showcasing preferential attachment, as neurons from the previous and next layers will connect only to these heads. The middle head (head $5$) is also maintained more than most other heads, as it replaces many of the earlier heads. Results averaged across GLUE tasks on the right are similar.
Redundancy-based pruning While structured sparsity aims at preferential attachment, such pruning (by zeroing out weights) cannot determine which connections are redundant. Elimination of redundant connections is an important aspect of the refinement process Wang et al. (2011); Hashimoto and Kano (2013) that takes place after the brain develops very dense networks of connections. In the case of LLMs, we hypothesize that such redundancy can be measured by similarity between attention heads, whereupon similar attention heads can then be merged, resulting in reduced complexity while maintaining functionality (i.e. performance is maintained on downstream tasks). Such removal of redundancy is conjectured in Lichtman and Colman (2020) to be unique to the neuron development in the central nervous system of vertebrates. Figure 3 illustrates how often heads are found to be redundant using NeuroPrune. Generally, the last head is found to be the least redundant, while the middle head also exhibits limited redundancy. In Figure 7(right), we see head removal as a function of layers and find that the last three layers have the highest number of redundant heads.
@@
Our approach has the following benefits:
  • It is task agnostic.
  • It is easily adaptable to different transformer-based LLM architectures as it does not add additional (mask) variables to do the pruning. We apply it to BERT (encoder), T5 (encoder-decoder) and OPT (decoder) models.
  • It learns sparsity patterns exhibiting principled topological structure.
  • It results in LLMs with a competitive and even sometimes superior performance on different benchmarks and tasks (GLUE, summarization, machine translation), although our proposal is more neuroscience-motivated than solely trying to maximize performance.
  • It is generally much faster to train than the competing baselines, with time per epoch being similar to standard fine-tuning. It also exhibits inference speedups as the topological constraints encourage $N$:$M$-type sparsity.
@@ 4 Method
Input: Model $M(\cdot;A_{0},L_{0})$ with $N_{L}$ layers, initial $Q$, $K$, $V$ concatenated matrices $A^{(l)}_{0}$ and MLP weight matrices $L^{(l)}_{0}$ for layer $l$, sparsity parameters $\alpha,\beta\geq 0$, # epochs $n_{e}$, loss function $\lambda$, redundancy threshold $\theta$
@@
Algorithm 1 NeuroPrune
@@ 4.1 MLP Sparsification
$R_{mlp}^{(l)}(L^{(l)})=$ $\frac{1}{d}[n^{(l)}_{in,1},...,n^{(l)}_{in,d_{e}}]\cdot|L_{in}^{(l)}|\cdot\vec{1}_{d}$ (1)
$+\frac{1}{d_{e}}[n^{(l)}_{out,1},...,n^{(l)}_{out,d}]\cdot|L_{out}^{(l)}|\cdot\vec{1}_{d_{e}}$
@@ 4.2 Attention Sparsification
Note that the above constraint is applied across heads in the attention layer as it considers the entire $Q$, $K$, $V$ matrices (hence the inner summation over $3d$ entries). Additionally, while the standard $l_{2}$ group penalty induces weights within a group to be similar, this $l_{1}^{.5}$ group penalty allows sparsity patterns to be learned within a group. For example, in Figures 1 and 8, while entire rows are often removed, we also observe that certain rows only exhibit sparsity in $Q$ and $K$ while leaving corresponding rows of $V$ dense, which is still valuable as it indicates that attention may not be required for those neurons/dimensions.
@@ 5 Experiments
Metrics: For performance we report accuracy (Acc) for the GLUE datasets except COLA where Mathew’s Correlation (MCorr) is a standard metric, Rouge for summarization and BLEU scores for machine translation. We also report (average) inference and train times. Please note that by train time we mean the total time taken to train the model using dynamic sparsity and inference time is the time taken by the model to process a test example based on the sparse model that has been learned.
Setup Details: NeuroPrune results were obtained by varying $\alpha$ and $\beta$ from $10^{-7}$ to $0.1$ in multiples of $10$. The $l_{1}$ penalty parameter also took these values. $\theta$ took values in $\{0.15,0.2,0.25\}$ for the head removal experiments where the default was set to $0.15$ for the GLUE and summarization tasks. It was sufficient to run NeuroPrune for a single epoch for the larger GLUE datasets (viz. MNLI, QQP, QNLI and SST2) and the summarization task, while for the other smaller GLUE datasets we ran it for $3$ to $5$ epochs. $\epsilon$ for MLP sparsification was set to $10^{-4}$. CoFI finetunes the model before it starts pruning. For smaller GLUE datasets, finetuning before pruning epochs is $4$ and total epochs is $100$, while for larger datasets, these numbers are $2$ and $20$ respectively. All the other parameters were unchanged. To make the comparison fair, we turned off the distillation option. DSP runs 3 epochs of joint (finetuning and mask learning) training. For machine translation, NeuroPrune results were obtained with $\alpha\in\{0.005,0.01,0.05,0.1,0.25,0.5\}$, $\beta\in\{0.05,0.1\}$ and $\theta$ varying from 0.1 to 0.4 in increments of 0.02. $\epsilon$ was set as in GLUE. Since there was no pretrained model, 15 epochs of pretraining were done followed by 15 epochs of NeuroPrune finetuning for a total of 30 epochs. DSP joint training was also done with a total of 30 epochs.
@@ 5.1 Varying sparsity percentage
GLUE: In Figure 4 we see the behavior of our method on GLUE datasets for BERT, T5, and OPT models. We see that our method is always competitive with CoFI, outperforming it on the smaller datasets. This is possible because we do not add extra variables to the model, which when coupled with the topological constraints results in stabler performance. The structured sparsity also gives improved inference times and the train time is much lower, since not only do we need to run our method only for a few epochs, but the time per epoch is the same as standard fine-tuning. CoFI on the other hand, as recommended in Xia et al. (2022), requires many epochs of running to reach specified sparsity levels as it first fine tunes and then prunes.
For T5 and OPT we compare with $l_{1}$ sparsification. As can be seen, NeuroPrune is better than $l_{1}$ in most cases w.r.t. performance, especially for in-between sparsities. We believe this happens because NeuroPrune can choose between attention or MLP ($\alpha$, $\beta$ parameters), on which to sparsify more when optimizing performance even though the constraints are structured. We also see benefits in inference time, again possibly because of the structured sparsification. Train times are similar as we run $l_{1}$ for the same number of epochs as NeuroPrune and its per-epoch time is similar to that of standard fine-tuning. Performance with varying sparsity on T5-large and OPT-1.3b are seen in Figure 11 in the appendix.
@@
Figure 6: Performance (left) and Efficient Frontier (right) for NeuroPrune and DSP on a German to English translation task. NeuroPrune consistently outperforms DSP across models with varying numbers of attention heads present and typically with much high levels of sparsity. The efficient frontier shows that for similar levels of performance and same number of heads, NeuroPrune finds much sparser models.
@@ 5.2 Varying number of heads
Figure 7: Left, we see the relative importance of removed heads for NeuroPrune and DSP averaged across runs, where at least 10 heads are removed. We see that NeuroPrune removes more important heads than DSP does because of its redundancy based head elimination algorithm. Right, we see the (average) # of heads removed per layer using NeuroPrune when fine tuning on the GLUE Benchmark datasets with a BERT(-base) model for cases where at least 10 heads are removed. As can be seen more heads are pruned from the later layers and the first one as compared to the middle layers.
@@ 5.3 Other Insights
Topological sparsity: As seen in Figure 1 (and Figures 8, 9, 10 in the appendix) we see that our constraints try to eliminate neurons both in the MLP layers as well as the attention layers. This is observed as a row/column (sparsity) pattern in these matrices. Interestingly, sometimes only the $Q$ and $K$ entries in a row are sparsified, although the group sparsity constraint is applied to $QKV$ jointly. The similarity in sparsification patterns for matrices $Q$ and $K$ directly reflects the symmetry of the operations they are jointly involved in, when computing the attention coefficients as inner products of each row of $Q$ with each row of $K$, compactly as $QK^{\top}$, unlike $V$ which is simply a linear projection of the token embeddings. The MLP portions, as seen in Figure 1, exhibit preferential attachment with increasing sparsity which is consistent with brain functional networks.
Importance of removed heads: In Figure 7(left), we see the relative importance of removed heads, defined as the ratio between the importance of removed heads in a layer to the ones that remain averaged across all the layers. The individual head importances are computed as the sum of the absolute output dense weights that emerge from a head. We see that NeuroPrune can prune more important heads than DSP as it finds redundant heads, thereby sparsifying more aggressively, which is evidenced by the faster training time on GLUE tasks for similar levels of head removal.
Head removal: As seen in Figure 5 there is a high bias to keep the last head in each layer. This is because NeuroPrune prefers keeping later heads in a layer if it is similar to a set of heads that another head may be similar to, thus encouraging more modularity in a layer. In Figure 7(right) in the appendix, we see that later layers, and often the first layer, lose more heads than the intermediate ones, which is consistent with Li et al. (2021).
@@ 6 Discussion
This work shows that NeuroPrune, inspired by sparsity, modularity, and preferential attachment topology in brain functional networks, is competitive and sometimes even superior to other SOTA dynamic sparse training methods, even though it does not solely try to optimize performance. It is also more efficient than them in train time with speed-ups also seen in inference. Not to mention it is easily transferable across transformer architectures (i.e. encoder, encoder-decoder, decoder) as it does not require adding additional gating variables to the architecture or modifying the architecture in any way. Also note that since dynamic sparse training and post-pruning are complimentary to each other one can potentially apply both, that is, the latter after the former. Moreover, since dynamic sparse training approaches employ sparsity during training the benefits of sparsity in terms of storage and also sometimes computation are leveraged both during training and inference. In contrast, post-pruning methods apply pruning after (dense) network training, benefiting from sparsity at inference alone.
There are multiple avenues for future research. First, it would be interesting to combine our redundancy-based and head-importance pruning methods to produce even more aggressive and efficient pruning of heads. Second, the structured strategies we used for fine-tuning could also be tried during pre-training. Third, one could test the generalizability of models trained using NeuroPrune on related, but different tasks and measure if similar gains can be secured. We believe our work will spur further progress on efficient dynamic sparse training methods that are also performant.
@@
All the datasets we considered were in English. Results may vary for other languages. We applied our method to three LLMs, but more architectures could be tested in the future not to mention more tasks. Our method although easy to adapt to different architectures while also being efficient, it does not allow the user to specify the exact number of heads one wants to prune and architectures are limited to Transformer based LLMs. This is implicitly a function of the threshold $\theta$ and the similarity of the pre-trained/fine tuned heads as well as the sparsity. We also have three hyper-parameters ($\alpha$, $\beta$ and $\theta$) that need to be specified for each run.
@@
Appendix A Additional Figures
In Figures 8, 9 and 10 we sparsity patterns induced by NeuroPrune in a BERT-base model at different sparsity percentages on the SST2 dataset.
Performance with varying sparsity on T5-large and OPT-1.3b are seen in Figure 11.
@@ Appendix A Additional Figures
[…]